# Rethinking Latent Representations in Behavior Cloning: An Information Bottleneck Approach for Robot Manipulation

### v2

@@ 5.3 Performance on LIBERO
Finding 5: For more complex language-conditioned multi-task scenarios, all baselines with different backbones incorporating IB consistently show performance improvements across all LIBERO benchmarks. For example, BC-VILT achieves large gains of 7.66% and 9.00% on LIBERO-Goal and LIBERO-Object, respectively, while BC-RNN shows a significant improvement of 10.83% on LIBERO-Goal. IB proves to be more effective in more complex environments and settings. We attribute this to the difference in task complexity: in CortexBench, the history length is 3, while in LIBERO, it is 10, with LIBERO being a multi-task benchmark and CortexBench being single-task benchmark. The increased data complexity (task quantity and input information) suggests a higher level of data redundancy, making IB even more effective.
Finding 6: We observe that in complex multi-task scenarios with more intricate inputs, such as a greater number of input modalities and extended historical information, using the Temporal Transformer in temporal fusion proves to be more effective than both spatial fusion and RNN-based temporal fusion. The evidence lies in the fact that the average success rates of BC-Transfomer and BC-VILT are over 30% higher than those of BC-MLP and BC-RNN. This is likely because Temporal Transformers excel in handling long-range interactions and capture dynamic dependencies across time steps, where RNNs and spatial fusion methods may struggle. This finding, together with Finding 1, underscores the specific scenarios in which each fusion method is most applicable.
Finding 7: IB is particularly effective for tasks requiring diverse feature extraction, such as distinguishing distinct task objectives or differentiating between various objects, as in LIBERO-Goal and LIBERO-Object. By filtering out irrelevant information, IB facilitates better generalization and more compact representations. However, its impact is less pronounced in spatial and long-horizon tasks, such as LIBERO-Spatial and LIBERO-Long, which heavily depend on structural and sequential dependencies that may be disrupted by excessive compression.
@@ 5.4 More Analysis
Effect of the Lagrange multiplier $\beta$ of Equation 6. This experiment evaluates how incorporating IB enhances performance. Since the MINE model’s parameters are fixed, the key difference between BC+IB and BC lies in the parameter $\beta$, which balances compression and predictive power. For LIBERO experiments, $\beta$ is explored within 1e-4, 1e-3, 5e-3, 1e-2. Finding 8: As shown in Figure 5, IB improves performance within a specific $\beta$ range, with a peak observed at an undetermined value. However, across all experiments, $\beta$ around 1e-4 consistently yields stable improvements. The selected $\beta$ values for each experiment are detailed in C.1.
Effect of the Number of Demonstrations. We evaluate IB’s effectiveness in few-shot settings, as few-shot learning is crucial for fine-tuning on domain-specific tasks in real-world applications. Finding 9: As shown in Figure 6, IB significantly improves performance even with limited data, highlighting its effectiveness in real-world scenarios where data is scarce. This further underscores the potential of IB in improving model generalization in practical settings.
Visualizations of $I(X,Z)$. As shown in Figure 4, BC+IB achieves a greater reduction in $I(X,Z)$ compared to vanilla BC, leading to significant performance improvements and further validating the effectiveness of IB. For instance, in LIBERO-Goal, incorporating IB reduces $I(X,Z)$ to one-quarter of its original value, resulting in a 7.7% increase in success rate.
We provide real-world experiments in Section C.2.
@@ 6 Limitations and Discussion
While our work provides extensive experimental validation of the effectiveness of IB and the necessity of input redundancy reduction in robotics representation learning, several limitations remain. First, for scalability, we do not incorporate large models like vision-language-action models due to high computational and time costs. Consequently, architectures like RT-2 (Brohan et al., 2023) and OpenVLA (Kim et al., 2024), which forgo a policy head and treat actions as text tokens, have not been explored. This is an avenue we aim to investigate in future work. Second, we have not examined alternative policy heads, such as diffusion-based (Chi et al., 2023) or transformer-based policy heads (Octo Model Team et al., 2024), which may further enhance performance. Third, while we evaluate our method on controlled benchmarks, its robustness to domain shifts, such as environmental or task variations, remains underexplored. We hope our work will inspire future research and contribute to the ongoing development and refinement of these methods.
@@ 7 Conclusions
In this study, we investigated the redundancy in representations for Behavior Cloning in robot manipulation and introduced the Information Bottleneck principle to mitigate this issue. By incorporating IB, we aimed to filter out redundant information in latent representations while preserving task-relevant features. Extensive experiments across various representation learning methods on CortexBench and LIBERO revealed insightful findings and demonstrated that IB significantly improves performance across diverse tasks and architectures. We hope our work will inspire further integration of information-theoretic principles into robotics and foster deeper theoretical analysis in this domain.
@@ B.1.1 CortexBench
We provide a detailed overview of the four imitation learning benchmarks used in CortexBench  (Majumdar et al., 2023). CortexBench is a single-task benchmark that includes 7 selected simulators, collectively offering 17 different embodied AI tasks spanning locomotion, navigation, and both dexterous and mobile manipulation. Three of the simulators are related to reinforcement learning, and thus are excluded from our analysis. The remaining four simulators, with a total of 14 tasks, are retained for validation: Adroit (2 tasks) (Rajeswaran et al., 2018), Meta-World (5 tasks) (Yu et al., 2020), DMControl (5 tasks) (Tassa et al., 2018), and TriFinger (2 tasks) (Wuthrich et al., 2021).
@@ B.3 Details of Implementations
For all experiments, we use a single Nvidia V100 GPU (CUDA 11.3) with 12 CPUs for training and evaluation.
@@ B.4.1 CortexBench
In the single-task dataset CortexBench, we observed that the learning curves of certain tasks exhibit significant oscillations, such as the assemble task in MetaWorld, as shown in Figure 9 a). Previous studies often record performance at intervals of many epochs or steps, selecting either the highest value (Majumdar et al., 2023) or the average of multiple peak values (Ze et al., 2024). Following (Majumdar et al., 2023), we directly use the highest value to explore the model’s full potential on the given task.
@@ B.4.2 LIBERO
For the multi-task dataset LIBERO, we observed that while the learning curve for individual tasks may still oscillate, a decrease in success rate for one task is often accompanied by an increase for another. This trade-off results in smoother overall learning curves across multiple tasks, as shown in Figure 9 b). To make model selection more practical and representative, we directly select the model from the final epoch.
@@ C.1 Details of Simulation Experiments
Figure 10: Real-world robot experiments conducted on a table setup involving two tasks. The left figure illustrates the experimental setup. The top right figure presents an example of the predicted particle trajectories alongside the policy execution. The bottom right figure provides the quantitative results,
C.2 Real-world Experiments
As shown in Figure 10, our real-world experiments involve a 6-DOF UR5 arm equipped with a Robotiq 2F-85 gripper and a RealSense L515 camera (base camera) for RGB image capture. We designed two simple tabletop manipulation tasks with manipulation skills:
  • Pick. The robot grips the pink cup on the table and lifts it up in the sky.
  • Pick and Place. The robot grips the pink cup on the table and places it in the bowl.
The demonstrations used for training both BC and BC+IB policies are collected using a 3D mouse, with 25 demonstrations recorded for pick tasks and 50 for more challenging pick-and-place tasks, utilizing only the base camera. We adopt VC-1 (Majumdar et al., 2023) as the baseline BC method, where pre-trained representations remain frozen during policy training, maintaining consistency with the simulation setup. To ensure a fair comparison, all methods are evaluated under identical initial conditions for each task.
The real-world robot experiments demonstrate that incorporating IB into BC significantly enhances task success rates. As shown in Table 10, BC+IB achieves a 9/10 success rate for picking the pink cup, compared to 5/10 for standard BC. Similarly, in the more complex task of placing the pink cup into the bowl, BC+IB outperforms BC with a success rate of 6/10 versus 3/10. This indicates that reducing redundancy in latent representations improves both grasping stability and overall task execution.
C.3 Extention to Few-shot Setting
We further evaluate the effectiveness of IB in a few-shot setting by conducting experiments with BC-VILT across multiple suites in LIBERO, as illustrated in Figure 11. The model is trained with only 10 demonstrations. The results consistently show that incorporating the IB improves success rates across all LIBERO suites, highlighting its efficacy in few-shot learning scenarios.
C.4 Experiments on LIBERO-Object of LIBERO
We observe that in the LIBERO-Object suite, the success rate does not consistently improve with an increasing number of demonstrations. Specifically, in the 10-shot setting, BC-VILT achieves a success rate of 56.17%, but in the full-shot setting, its performance drops to 43.00%. We hypothesize that this decline stems from inherent data distribution characteristics within the benchmark. In particular, we suspect that task-level imbalance in the dataset may contribute to overfitting.

### v3

@@ 5.3 Performance on LIBERO
Finding 6: For more complex language-conditioned multi-task scenarios, all baselines with different backbones incorporating IB consistently show performance improvements across all LIBERO benchmarks. For example, BC-VILT achieves large gains of 7.66% and 9.00% on LIBERO-Goal and LIBERO-Object, respectively, while BC-RNN shows a significant improvement of 10.83% on LIBERO-Goal. IB proves to be more effective in more complex environments and settings. We attribute this to the difference in task complexity: in CortexBench, the history length is 3, while in LIBERO, it is 10, with LIBERO being a multi-task benchmark and CortexBench being a single-task benchmark. The increased data complexity (task quantity and input information) suggests a higher level of data redundancy, making IB even more effective.
Finding 7: We observe that in complex multi-task scenarios with more intricate inputs, such as a greater number of input modalities and extended historical information, using the Temporal Transformer in temporal fusion proves to be more effective than both spatial fusion and RNN-based temporal fusion. The evidence lies in the fact that the average success rates of BC-Transformer and BC-VILT are over 30% higher than those of BC-MLP and BC-RNN. This is likely because Temporal Transformers excel in handling long-range interactions and capturing dynamic dependencies across time steps, where RNNs and spatial fusion methods may struggle. This finding, together with Finding 1, underscores the specific scenarios in which each fusion method is most applicable.
Finding 8: IB is particularly effective for tasks requiring diverse feature extraction, such as distinguishing distinct task objectives or differentiating between various objects, as in LIBERO-Goal and LIBERO-Object. By filtering out irrelevant information, IB facilitates better generalization and more compact representations. However, the impact is less pronounced in spatial tasks such as LIBERO-Spatial, which rely heavily on structural information that can be disrupted by excessive compression. For long-horizon tasks, the main performance bottleneck lies in the lightweight baseline models, whose limited capacity restricts their effectiveness on more complex tasks such as LIBERO-Long. This limitation is further illustrated and solved in Appendix Table 5.
5.4 Performance on Real World Experiments
As shown in Figure 4 (b) and (c), incorporating IB consistently improves success rates across both single-task and language-conditioned multi-task real-world settings in most cases. In the single-task setting, VC1+IB significantly outperforms VC1 in both the pick and put tasks. In the more challenging language-conditioned multi-task setting, CogAct+IB consistently outperforms CogAct across most tasks, including unseen object–bowl combinations, demonstrating enhanced generalization capabilities. These results suggest that reducing redundancy in latent representations leads to more robust grasping and more reliable overall execution in real-world scenarios.
@@ 5.5 More Analysis
Effect of the Lagrange multiplier $\beta$ of Equation 6. This experiment evaluates how incorporating IB enhances performance. Since the MINE model’s parameters are fixed, the key difference between BC+IB and BC lies in the parameter $\beta$, which balances compression and predictive power. For LIBERO experiments, $\beta$ is explored within 1e-4, 1e-3, 5e-3, 1e-2. Finding 9: As shown in Figure 5, IB improves performance within a specific $\beta$ range, with a peak observed at an undetermined value. However, across all experiments, $\beta$ around 1e-4 consistently yields stable improvements. The selected $\beta$ values for each experiment are detailed in C.1.
Effect of the Number of Demonstrations. We evaluate IB’s effectiveness in few-shot settings, as few-shot learning is crucial for fine-tuning on domain-specific tasks in real-world applications. As shown in Figure 7, IB consistently improves performance even with limited data across multiple suites in LIBERO, highlighting its effectiveness in real-world scenarios where data is scarce. This further underscores the potential of IB in improving model generalization in practical settings.
Visualizations of $I(X,Z)$. As shown in Figure 6, BC+IB achieves a larger reduction in $I(X,Z)$ compared to vanilla BC, leading to improved performance and validating the effectiveness of IB. For example, in LIBERO-Goal, IB reduces $I(X,Z)$ to one-quarter of its original value and yields a 7.7% increase in success rate.
@@ 6 Limitations and Discussion
While our work provides extensive experimental validation of the effectiveness of IB and the necessity of input redundancy reduction in robotics representation learning, several limitations remain. First, we do not comprehensively investigate the scalability of IB in robotic models. Most of our experiments are conducted on relatively lightweight architectures, with only the real-world experiments utilizing CogAct. We have not systematically evaluated large models such as vision-language-action architectures, due to the high computational and time costs involved. Second, alternative policy heads such as transformer-based designs, and models like RT-2 (Brohan et al., 2023) and OpenVLA  (Kim et al., 2024), which remove the explicit policy head and instead treat actions as text tokens, have not been explored. We leave these directions for future work. Third, although we evaluate our method on various benchmarks, its robustness to domain shifts, including variations in environment and task settings, remains underexplored. We hope our work will inspire future research and contribute to the ongoing development and refinement of these methods.
@@ 7 Conclusions
In this study, we investigated the redundancy in representations for Behavior Cloning in robot manipulation and introduced the Information Bottleneck principle to mitigate this issue. By incorporating IB, we aimed to filter out redundant information in latent representations while preserving task-relevant features. Extensive experiments across various representation learning methods on CortexBench and LIBERO revealed insightful findings and demonstrated that IB consistently improves performance across diverse tasks and architectures. We hope our work will inspire further integration of information-theoretic principles into robotics and foster deeper theoretical analysis in this domain.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning in Robotics. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.1.1 CortexBench
We provide a detailed overview of the four imitation learning benchmarks used in CortexBench (Majumdar et al., 2023). CortexBench is a single-task benchmark that includes 7 selected simulators, collectively offering 17 different embodied AI tasks spanning locomotion, navigation, and both dexterous and mobile manipulation. Three of the simulators are primarily designed for reinforcement learning and are therefore excluded from our analysis. The remaining four simulators, with a total of 14 tasks, are retained for validation: Adroit (2 tasks) (Rajeswaran et al., 2018), Meta-World (5 tasks) (Yu et al., 2020), DMControl (5 tasks) (Tassa et al., 2018), and TriFinger (2 tasks) (Wuthrich et al., 2021).
@@ B.3 Details of Implementations
For Cortexbench and LIBERO experiments, we use a single NVIDIA V100 or A100 GPU (CUDA 11.8) with 12 CPUs. For real-world experiments, the single-task setting uses one V100 GPU with 12 CPUs, while the language-conditioned multi-task setting is trained on 8 A100 GPUs with 100 CPUs and evaluated on a single A100 GPU.
@@ B.4.1 CortexBench
In the single-task dataset CortexBench, we observed that the learning curves of certain tasks exhibit significant oscillations, such as the assemble task in MetaWorld, as shown in Figure 10 (a). Previous studies often record performance at intervals of many epochs or steps, selecting either the highest value (Majumdar et al., 2023) or the average of multiple peak values (Ze et al., 2024). Following (Majumdar et al., 2023), we directly use the highest value to explore the model’s full potential on the given task.
@@ B.4.2 LIBERO
For the multi-task dataset LIBERO, we observed that while the learning curve for individual tasks may still oscillate, a decrease in success rate for one task is often accompanied by an increase for another. This trade-off results in smoother overall learning curves across multiple tasks, as shown in Figure 10 (b). To make model selection more practical and representative, we directly select the model from the final epoch.
@@ C.1 Details of Simulation Experiments
Figure 11: IB encourages the model to focus more selectively on task-relevant regions, specifically the robotic arm and the target object, by suppressing attention to redundant background features.
C.2 Attention Map Visualizations
As shown in Figure 11, IB helps the model focus on task-relevant regions (e.g., the arm and target object) by suppressing attention to redundant background features—an effect less evident without redundancy reduction, highlighting IB’s distinct role beyond standard regularization.
C.3 More Explanation on the LIBERO Experiments
C.3.1 More Experiments on LIBERO-Long
In the main text, we point out that the experimental results on LIBERO-Long are affected by the limited capacity of baseline models (e.g., BC-VILT with 10M parameters), leading to only marginal improvements. While these models perform well on simpler tasks such as LIBERO-Goal, achieving an 80% success rate with an 8% improvement, they show limited gains on more complex tasks like LIBERO-Long, with only a 2% increase in performance. This is primarily due to the performance ceiling imposed by the lightweight baselines. To further illustrate this limitation, we conduct experiments on LIBERO-Long using Diffusion Policy (DP), where the 1.14M-parameter MLP head of BC-Transformer is replaced with the more expressive 90M-parameter DP head. We train DP from scratch on a single A100 GPU with a batch size of 64, a learning rate of 1e-4, and for 50 epochs. The training setup for DP+IB is identical to that of DP, except for the IB-related components. For the IB-specific hyperparameters, the Lagrange multiplier $\beta$ is set to 1e-5.
Table 10: Performance on LIBERO-Long.
Method LIBERO-Long
DP 78.0
DP+IB 84.0
Attention: Our experimental setup differs from that of papers like OpenVLA. In those works, the image observations are saved at a resolution of 256×256 (instead of 128×128) and undergo additional filtering, such as removing ”no-op” (zero) actions and unsuccessful demonstrations. In contrast, our setting uses the raw LIBERO data with lower-resolution images and no filtering.
C.3.2 Performance on LIBERO-Object
On LIBERO-Object, we observe that the success rate does not consistently improve with an increasing number of demonstrations. Specifically, in the 10-shot setting, BC-VILT achieves a success rate of 56.17%, whereas in the full-shot setting, its performance drops to 43.00%. We hypothesize that this decline is due to inherent data distribution characteristics and suboptimal data quality within the benchmark.
C.4 Extension to Few-shot Setting
We further assess the effectiveness of IB in few-shot settings by evaluating BC-VILT under varying numbers of demonstrations in LIBERO-Goal, as shown in Figure 12. The results consistently show that incorporating IB improves success rates across all LIBERO suites, demonstrating its efficacy in few-shot learning scenarios.